OpenStudy (anonymous):

A person plays a game where he blindly picks a colored chip from a container. There are 20 chips in the container: 5 red, 11 green, and 4 blue. He gets $1 for a red chip, nothing for a green chip, and he has to pay $1.25 if he picks a blue chip. This is a fair game. True False

OpenStudy (welshfella):

work out the probabilities of picking a blue chip and a red chip

OpenStudy (anonymous):

red chip=5/20=0.25 blue chip=4/20= 0.2

OpenStudy (welshfella):

on the face of it that seems fair as there is less chance of getting a blue than a red, The money is in the same ratio as the probabilities. But the chaance of picking a green is greater than 1/2...

OpenStudy (welshfella):

so what do you think?

OpenStudy (welshfella):

- but he doesn't have to pay money to play the game.

OpenStudy (anonymous):

okay that sounds fair to me @welshfella

OpenStudy (welshfella):

yes i think it is fair If you hade to pay to play then it would not be fair